# Preferential Latent Space Models for Networks with Textual Edges

### v2

@@ 3 Estimation
Given network data $Y=\{{Y}_{ij}\}_{i,j\in[n]}$, where ${Y}_{ij}\in[0,1]^{m_{ij}\times K}$, we aim to estimate ${a}$, ${W}$, and ${U}$. Under model (1), the negative log-likelihood function, up to a constant, can be written as
$\ell\left(a,{W},{U}\right)=-\sum_{k=1}^{K}\sum_{i<j}^{n}\sum_{l=1}^{m_{ij}}\frac{1}{m_{ij}}\left\{y^{(k)}_{ijl}\Theta^{(k)}_{ij}+\log\left(1-\psi(\Theta^{(k)}_{ij})\right)\right\},$ (2)
@@
where $s$ is a tuning parameter that controls the sparsity of ${W}$. To enforce sparsity and the positivity constraint on ${W}$ along the solution path, we employ a truncation operator Truncate(${W},s$) for $s\leq nK$ such that $[\operatorname{Truncate}({W},s)]_{ik}=W_{ik}$ if $(i,k)\in\text{Supp}^{+}({W},s)$ and $[\operatorname{Truncate}({W},s)]_{ik}=0$ otherwise, where $\text{Supp}^{+}({W},s)$ denotes the $s$ entries in ${W}$ with the largest values. To solve (3), we develop a projected gradient descent algorithm that is easy to implement and computationally efficient. Our estimation procedure is summarized in Algorithm 1.
@@
  Input: network data $Y$, initial values ${a}^{(0)},\bm{W}^{(0)},{U}^{(0)}$, step sizes $\eta_{{a}},\eta_{{W}},\eta_{{U}}$.
@@
The parameters $\eta_{{a}},\eta_{{W}},\eta_{{U}}$ control the step sizes in the gradient descent algorithm. Theorem 1 provides theoretical conditions on $\eta_{{a}},\eta_{{W}},\eta_{{U}}$ to ensure the algorithm achieves a linear convergence rate. In practice, backtracking line search can be implemented for $\eta_{{a}},\eta_{{W}},\eta_{{U}}$ at each step of the iteration to achieve fast convergence.
For the initialization of Algorithm 1, we consider a singular value thresholding based approach (Ma et al., 2020), which has demonstrated good empirical performance. See Section S1 in the supplement for details. The latent dimension $d$ and the sparsity $s$ are two tuning parameters in the proposed model. We select these two parameters using edge cross-validation (Li et al., 2020). Specifically, we divide all indices $\{i,j,l,k\}$’s into $L$ folds and use each fold as a validation set while training the model on the remaining $L-1$ folds. To calculate the cross-validation error on the validation set, we calculate the binomial deviance, and the $d$ and $s$ combination with the smallest cross-validation error is selected.
@@ 4 Theoretical Results
$\mathcal{F}_{n,d,K}(M_{1})=\bigg{\{}({a},{W},{U})\mid$ $\,\,\|{a}\|_{\infty}\leq{M_{1}}/{4},\,\,\max_{i}\sum_{k=1}^{K}(W_{ik})^{2}\leq{M_{1}}/{2},\,\,\|{W}\|_{0}<nK,$ (4)
$\|u_{i}\|_{2}=1,\,\,\max_{i\neq j,k}\Theta_{ij}^{(k)}\leq-(1-C)M_{1}\bigg{\}},$
where $M_{1}\geq 0$ is a scalar that may depend on $n$ and $10<C<$ 1 is a constant. By the definition of $\Theta_{ij}^{(k)}$ in (1) and combining $\|{a}\|_{\infty}\leq{M_{1}}/{4}$, $\max_{i}\sum_{k=1}^{K}(W_{ik})^{2}\leq{M_{1}}/{2}$ and $\|u_{i}\|_{2}=1$ in (4), it is straightforward to show that $\max_{i,j,k}|\Theta_{ij}^{(k)}|\leq M_{1}$. Hence, for any $({a},{W},{U})\in\mathcal{F}_{n,d,K}(M_{1})$, $\Theta_{ij}^{(k)}$’s are uniformly bounded in $[-M_{1},-(1-C)M_{1}]$ for any $k$ and $i\neq j$. That is, edge probabilities $\psi({\Theta_{ij}^{(k)}})$’s are bounded between ${1}/({1+e^{M_{1}}})$ and ${1}/({1+e^{(1-C)M_{1}}})$. It is seen that $M_{1}$ controls the overall sparsity of the network. If, for example, $M_{1}$ is in the order of $\log(n)-\log\log(n)$, then the average edge probability is in the order of $\log(n)/n$.
@@
We first introduce an error metric for the iterates from Algorithm 1. As ${U}$ is identifiable up to an orthogonal transformation, for any ${U}_{1},{U}_{2}\in\mathbb{R}^{n\times d}$, we define a distance measure
$\text{dist}({U}_{1},{U}_{2})=\min_{{R}:{R}{R}^{\top}={I}}\|{U}_{1}-{U}_{2}{R}\|_{F}.$
@@
This assumption requires the initial values to be reasonably close to the true parameters. Such assumptions are commonly employed in nonconvex optimizations (Lyu et al., 2023; Zhang et al., 2023). In particular, if $\kappa_{0}=O(1)$ and $d=O(1)$, then Assumption 1 can be simplified to $\|{a}^{(0)}-{a}^{\ast}\|_{2}^{2}=O(nw_{\max}^{4}e^{-2M_{1}})$, $\|{W}^{(0)}-{W}^{\ast}\|_{F}^{2}=O(Knw_{\max}^{2}e^{-2M_{1}})$ and $\text{dist}^{2}({U}^{(0)},{U}^{\ast})=O(ne^{-2M_{1}})$. These assumptions on ${a}^{(0)}$, ${W}^{(0)}$ and ${U}^{(0)}$ are relatively mild.
@@
Suppose $({a}^{(0)},{W}^{(0)},{U}^{(0)})$ satisfies Assumption 1 and $({a}^{\ast},{W}^{\ast},{U}^{\ast})$ is in (4) and satisfies Assumption 2. Letting $\eta_{a}={\eta}/({4Kn})$, $\eta_{{W}}={\eta}/({4{\sigma_{1}^{\ast}}^{2}w_{\max}^{2}})$, $\eta_{{U}}={\eta}/({2K{\sigma_{1}^{\ast}}^{2}w_{\max}^{4}})$ and $s=\gamma s^{\ast}$ for $\gamma>1$, the $t$-th step iteration of Algorithm 1 satisfies, with probability as least $1-Kn^{-1}$,
@@
Our theoretical analysis is nontrivial as the node-layer preferential effects in $W_{ik}W_{jk}u_{i}u_{j}$ lead to an involved interplay between ${W}$ and ${U}$. This requires carefully bounding the error of ${W}^{(t)}$ and ${U}^{(t)}$ (up to rotation) in each step of the iteration to achieve contraction while ensuring the identifiability conditions are met. To tackle varying edge number $m_{ij}$’s, we derive a tight bound on the spectrum of random matrices with bounded moments following the techniques in Bandeira and van Handel (2016); see Lemma S5. This bound is sharper than the matrix Bernstein inequality (Tropp, 2012). Using Lemma S5, we are able to improve the statistical error for low rank matrix ${U}^{\ast}$ from $\kappa_{0}^{4}e^{(1+C)M_{1}}dn\log(n)$, which can be derived using the matrix Bernstein inequality under $m_{ij}=1$, to $\kappa_{0}^{4}e^{(1+C)M_{1}}d\max\left\{{n}/{\bar{m}},\log(n)\right\}$, which in turn relaxes the minimal signal strength condition in Assumption 2.
@@
In Corollary 1, the error bounds for ${a}^{(t)}$ and ${U}^{(t)}$ decrease with $K$, indicating that their estimation improves as the number of layers $K$ increases. All three estimation errors decrease with the average number of edges $\bar{m}$, suggesting that observing more edges between nodes leads to better estimation. Finally, the estimation error for ${U}^{\ast}$ matches with that in standard latent space models (Ma et al., 2020) when $K=1$, $\bar{m}=1$ and $W_{ij}=1$ for all $i,j$.
@@ 5 Simulation
$\frac{\|\hat{{a}}-{a}^{\ast}\|_{2}^{2}}{\|{a}^{\ast}\|^{2}_{2}},\,\,\frac{\|\hat{W}-W^{\ast}\|_{F}^{2}}{\|W^{\ast}\|_{F}^{2}},\,\,\min_{R:\,{R}^{\top}{R}={R}{R}^{\top}={I}_{k}}\frac{\left\|\hat{{U}}-{U}^{\ast}R\right\|_{F}^{2}}{\left\|{U}^{\ast}\right\|_{F}^{2}},\,\,\frac{1}{K}\sum_{k=1}^{K}\frac{\left\|\psi(\hat{\Theta}^{(k)})-\psi({\Theta}^{(k)}{}^{\ast})\right\|_{F}^{2}}{\left\|\psi({{{\Theta}^{(k)}}{}^{\ast}})\right\|_{F}^{2}}.$
where $\hat{{a}}$, $\hat{{W}}$ and $\hat{{U}}$ are estimators from Algorithm 1, and $\psi({\Theta}^{(k)}{}^{\ast})$ is true edge probability calculated using ${a}^{\ast}$, ${W}^{\ast}$ and ${U}^{\ast}$.
@@
Figures 2-3 report the estimation errors of ${{a}}^{\ast}$, ${{W}}^{\ast}$, ${U}^{\ast}$ and $\psi({\Theta}^{(k)}{}^{\ast})$ under various settings, with 95% confidence intervals, over 100 data replications.
Figure 3: Mean relative errors and their corresponding $95\%$ intervals under varying edge density while $n=200$, $K=20$ and $m=1$ (top panel), and under varying $m$ while $n=200$, $K=20$ and edge density at 0.08 (bottom panel).
@@ 6.1 Data description
The Enron email corpus (Klimt and Yang, 2004), one of the most extensive publicly available datasets of its kind, contains over 500,000 emails from 158 employees from November 13, 1998 to June 21, 2002. Released by the Federal Energy Regulatory Commission following its investigation of Enron, this dataset provides a unique opportunity to study the communications and organizations within a major corporation during various stages of a financial collapse. The study period can be divided into three stages, as marked by two major events. In February 2001, Enron’s stock reached its peak and then began a dramatic decline following major sell-offs from top executives. It was later found that starting February 2001, concerns about Enron’s accounting practices were increasingly discussed internally. In October 2001, the company’s financial scandal was publicly exposed and the Securities and Exchange Commission began an investigation into Enron’s accounting practices. Accordingly, we consider three stages in our analysis: the pre-decline period from November, 1998, to February, 2001; the decline and pre-bankruptcy period from February, 2001, to October, 2001; and the bankruptcy and post-bankruptcy period from October, 2001, to June, 2002.
Keywords
Enron, corp, gas, power, energy, market, agreement,
information, price, deal, contract, credit, legal, fax, review,
capacity, phone, north, america,draft, texas, buy, trade,
discuss, risk, issue, position, letter, plan, stock,
natural, sell, value, customer, pipeline, product, continue,
cash, physical, pay, plant, account, picture, attach, earn
Table 3: Extracted keywords sorted by their frequency of occurrences in the corpus.
Our analysis focused on the emails of 154 employees whose roles and departments are documented in the dataset. These employees are from different departments including gas, legal, transportation, regulatory and government, and wholesale, among others. Data preparation included standard text preprocessing steps such as tokenization, lower-casing, removel of punctuations and stopwords, and stemming. We then employed latent Dirichlet allocation to extract keywords that inform latent topics in the corpus. The 45 extracted keywords are shown in Table 3. To construct the network for each stage, we merged all emails between each pair of nodes into a single document. This process created an undirected network with $n=154$ nodes and $K=45$ keywords for each stage, denoted by $\{y_{ij}^{(k)}\}_{i,j\in[154],k\in[45]}$, where $y_{ij}^{(k)}=y_{ji}^{(k)}=1$ if the merged document between users $i$ and $j$ includes keyword $k$, and $y_{ij}^{(k)}=y_{ji}^{(k)}=0$ otherwise. This approach helps to reduce the sparsity of the network and facilitates comparison with other multi-layer network analysis methods. The edge densities of the three networks constructed for stages 1, 2, and 3 are 0.012, 0.010, and 0.009, respectively.

### v3

@@ 3 Estimation
Given network data $Y=\{{Y}_{ij}\}_{i,j\in[n]}$, where ${Y}_{ij}\in[0,1]^{m_{ij}\times K}$, we aim to estimate model parameters ${a}$, ${W}$, and ${U}$. Under model (1), the negative log-likelihood function, up to a constant, can be written as
$\ell\left(a,{W},{U}\right)=-\sum_{k=1}^{K}\sum_{i\leq j}^{n}\sum_{l=1}^{m_{ij}}\frac{1}{m_{ij}}\left\{y^{(k)}_{ijl}\Lambda^{(k)}_{ij}+\log\left(1-\psi(\Lambda^{(k)}_{ij})\right)\right\},$ (2)
@@
where $s$ is a tuning parameter that controls the sparsity of ${W}$. To enforce sparsity and the positivity constraint on ${W}$ along the solution path, we employ a truncation operator Truncate(${W},s$) defined as,
$[\operatorname{Truncate}({W},s)]_{ik}=\begin{cases}W_{ik},&\text{ if }(i,k)\in\text{Supp}^{+}({W},s),\\ 0,&\text{ otherwise, }\end{cases}$
for ${W}\in\mathbb{R}^{n\times K}$ and $s<nK$. The set $\text{Supp}^{+}({W},s)$ denotes the $s$ entries in ${W}$ with the largest values. To solve (3), we consider a projected gradient descent algorithm that is easy to implement and computationally efficient. Our estimation procedure is summarized in Algorithm 1.
@@
  Input: network data $Y$, initial values ${a}^{(0)},W^{(0)},{U}^{(0)}$, step sizes $\eta_{{a}},\eta_{{W}},\eta_{{U}}$.
@@
The parameters $\eta_{{a}},\eta_{{W}},\eta_{{U}}$ control the step sizes in the gradient descent algorithm. Theorem 1 provides theoretical conditions on $\eta_{{a}},\eta_{{W}},\eta_{{U}}$ to ensure the algorithm achieves a linear convergence rate. In practice, backtracking line search can be implemented for $\eta_{{a}},\eta_{{W}},\eta_{{U}}$ at each step of the iteration to achieve fast convergence. For the initialization of Algorithm 1, we consider a singular value thresholding based approach (Ma et al., 2020), which has demonstrated good empirical performance. See Section S1 in the supplement for details.
The latent dimension $d$ and the sparsity $s$ are two tuning parameters in the proposed model. We select these two parameters using edge cross-validation. Specifically, we divide all indices $\{i,j,l,k\}$’s into $L$ folds and use each fold as a validation set while training the model on the remaining $L-1$ folds. To calculate the cross-validation error on the validation set, we consider the binomial deviance, and the $d$ and $s$ combination with the smallest cross-validation error is selected.
@@ 4 Theoretical Results
$\Omega_{n,d,K}(M_{1})=\bigg{\{}({a},{W},{U})\mid$ $\,\,\|{a}\|_{\infty}\leq{M_{1}}/{4},\,\,\max_{i}\sum_{k=1}^{K}(W_{ik})^{2}\leq{M_{1}}/{2},\,\,\|{W}\|_{0}<nK,$ (4)
$\|u_{i}\|_{2}=1,\,\,\max_{i,j,k}\Lambda_{ij}^{(k)}\leq-(1-C)M_{1}\bigg{\}},$
where $M_{1}\geq 0$ is a scalar that may depend on $n$ and $10<C<$ 1 is a constant. By the definition of $\Lambda_{ij}^{(k)}$ in (1) and combining $\|{a}\|_{\infty}\leq{M_{1}}/{4}$, $\max_{i}\sum_{k=1}^{K}(W_{ik})^{2}\leq{M_{1}}/{2}$ and $\|u_{i}\|_{2}=1$ in (4), it is straightforward to show that $\max_{i,j,k}|\Lambda_{ij}^{(k)}|\leq M_{1}$. Hence, for any $({a},{W},{U})\in\Omega_{n,d,K}(M_{1})$, $\Lambda_{ij}^{(k)}$’s are uniformly bounded in $[-M_{1},-(1-C)M_{1}]$ for any $i,j$ and $k$. That is, edge probabilities $\psi({\Lambda_{ij}^{(k)}})$’s are bounded between ${1}/({1+e^{M_{1}}})$ and ${1}/({1+e^{(1-C)M_{1}}})$. It is seen that $M_{1}$ controls the overall sparsity of the network. If, for example, $M_{1}$ is in the order of $\log(n)-\log\log(n)$, then the average edge probability is in the order of $\log(n)/n$.
@@
To investigate the computational and statistical properties of iterates from Algorithm 1, we first introduce an error metric for the iterates from Algorithm 1. As ${U}$ is identifiable up to an orthogonal transformation, for any ${U}_{1},{U}_{2}\in\mathbb{R}^{n\times d}$, we define a distance measure
$\text{dist}({U}_{1},{U}_{2})=\min_{{R}:{R}{R}^{\top}={I}_{d}}\|{U}_{1}-{U}_{2}{R}\|_{F}.$
@@
This assumption requires the initial values to be reasonably close to the true parameters. Such assumptions are commonly employed in nonconvex optimizations (Lyu et al., 2023; Zhang et al., 2023). In particular, if $\kappa_{0}=O(1)$ and $d=O(1)$, then Assumption 1 can be simplified to $\|{a}^{(0)}-{a}^{\ast}\|_{2}^{2}=O(nw_{\max}^{4}e^{-2M_{1}})$, $\|{W}^{(0)}-{W}^{\ast}\|_{F}^{2}=O(Knw_{\max}^{2}e^{-2M_{1}})$ and $\text{dist}^{2}({U}^{(0)},{U}^{\ast})=O(ne^{-2M_{1}})$. These assumptions on ${a}^{(0)}$, ${W}^{(0)}$ and ${U}^{(0)}$ are mild.
@@
Suppose $({a}^{(0)},{W}^{(0)},{U}^{(0)})$ satisfies Assumption 1, $({a}^{\ast},{W}^{\ast},{U}^{\ast})$ is in (4) and satisfies Assumption 2, and $d\kappa_{0}^{8}e^{3M_{1}}=O({\sigma_{1}^{\ast}}^{2})$. Letting $\eta_{a}={\eta}/({4Kn})$, $\eta_{{W}}={\eta}/({4{\sigma_{1}^{\ast}}^{2}w_{\max}^{2}})$, $\eta_{{U}}={\eta}/({2K{\sigma_{1}^{\ast}}^{2}w_{\max}^{4}})$ and $s=\gamma s^{\ast}$ for $\gamma>1$, the $t$-th step iteration of Algorithm 1 satisfies, with probability as least $1-Kn^{-1}$,
@@
Compared with other work on network latent space models (Ma et al., 2020; Zhang et al., 2020), our theoretical analysis faces a few unique challenges. First, the node-topic preferential effects in $W_{ik}W_{jk}u_{i}u_{j}$ lead to an involved interplay between ${W}$ and ${U}$, as these effects vary across different topics and node pairs. This requires carefully bounding the error of ${W}^{(t)}$ and ${U}^{(t)}$ (up to rotation) separately in each step of the iteration to achieve contraction while ensuring the identifiability conditions are met. Second, the edge number $m_{ij}$’s vary across node pairs. To tackle varying edge numbers, we derive a tight bound on the spectrum of random matrices with bounded moments following the techniques in Bandeira and van Handel (2016); see Lemma S5. The proof of this result involves intricate technical details, and it uses large deviation estimates and geometric functional analysis techniques. The resulting bound is sharper than the matrix Bernstein inequality (Tropp, 2012). Using Lemma S5, we are able to improve the statistical error for low rank matrix ${U}^{\ast}$ from $\kappa_{0}^{4}e^{(1+C)M_{1}}dn\log(n)$, which can be derived using the matrix Bernstein inequality under $m_{ij}=1$, to $\kappa_{0}^{4}e^{(1+C)M_{1}}d\max\left\{{n}/{\bar{m}},\log(n)\right\}$, which in turn relaxes the minimal signal strength condition in Assumption 2. Lemma S5 extends the result in Lei and Rinaldo (2015), which was derived for the case of $m_{ij}=1$ using a different technique and highlights the benefit of having a greater average number of edges $\bar{m}$. Finally, the theoretical analysis is nontrivial, as it involves alternating gradient descent, orthogonal transformation, identifiability constraints, sparsity, and a non-quadratic loss function.
Based on Theorem 1, we can further derive the following error bounds for the estimated model parameters.
@@
In Corollary 1, the error bounds for ${a}^{(t)}$ and ${U}^{(t)}$ decrease with $K$, indicating that their estimation improves as the number of layers $K$ increases. The error bound for $W$ does not improve with $K$, as $W$ is not a common parameter shared across layers. All three estimation errors decrease with the average number of edges $\bar{m}$, suggesting that observing more edges between nodes leads to better estimation. Finally, the estimation error for ${U}^{\ast}$ matches with that in standard latent space models (Ma et al., 2020) when $K=1$, $\bar{m}=1$ and $W_{ij}=1$ for all $i,j$.
@@ 5 Simulation
$\frac{\|\hat{{a}}-{a}^{\ast}\|_{2}^{2}}{\|{a}^{\ast}\|^{2}_{2}},\,\,\frac{\|\hat{W}-W^{\ast}\|_{F}^{2}}{\|W^{\ast}\|_{F}^{2}},\,\,\min_{{R}:\,{R}^{\top}{R}={R}{R}^{\top}={I}_{k}}\frac{\left\|\hat{{U}}-{U}^{\ast}{R}\right\|_{F}^{2}}{\left\|{U}^{\ast}\right\|_{F}^{2}},$
where $\hat{{a}}$, $\hat{{W}}$ and $\hat{{U}}$ denote the estimators from Algorithm 1. Also reported is the relative estimation error of edge probabilities $\Lambda$, calculated as
$\frac{1}{K}\sum_{k=1}^{K}\frac{\left\|\psi(\hat{\Lambda}^{(k)})-\psi({\Lambda}^{(k)^{\ast}})\right\|_{F}^{2}}{\left\|\psi({{\Lambda^{(k)}}^{\ast}})\right\|_{F}^{2}},$
where ${\Lambda^{(k)}}^{\ast}$ is true edge probability calculated using ${a}^{\ast}$, ${W}^{\ast}$ and ${U}^{\ast}$.
@@
Figures 2-3 report the estimation errors of ${{a}}^{\ast}$, ${{W}}^{\ast}$, ${U}^{\ast}$ and $\psi({\Lambda}^{(k)}{}^{\ast})$ under various settings, with 95% confidence intervals, over 100 data replications. We apply the cross-validation procedure described in Section 3 to select the latent dimension $d$, and it consistently identifies the correct value of $d=2$.
Figure 3: Mean relative errors and their corresponding $95\%$ intervals under varying edge density while $n=200$, $K=20$, and $m=1$ (top panel), and under varying $m$ while $n=200$, $K=20$ and edge density at 0.08 (bottom panel).
@@ 6.1 Data description
The Enron email corpus (Klimt and Yang, 2004), one of the most extensive publicly available datasets of its kind, contains over 500,000 emails from 158 employees from November 13, 1998 to June 21, 2002. This dataset is released by the Federal Energy Regulatory Commission following its investigation of Enron. By analyzing this dataset with our proposed method, we offer an enriched view of the communications during one of the largest bankruptcy reorganizations in the U.S. history.
@@
The study period can be divided into three stages, as marked by two major events. In February 2001, Enron’s stock reached its peak and then began a dramatic decline following major sell-offs from top executives. It was later found that starting February 2001, concerns about Enron’s accounting practices were increasingly discussed internally. In October 2001, the company’s financial scandal was publicly exposed and the Securities and Exchange Commission began an investigation into Enron’s accounting practices. Accordingly, we consider three stages in our analysis: the pre-decline period from November, 1998 to February, 2001; the decline and pre-bankruptcy period from February, 2001 to October, 2001; and the bankruptcy and post-bankruptcy period from October, 2001 to June, 2002.
Table 3: Extracted topics from the Enron email corpus. Within each category, topics are sorted in descending order based on their frequency of occurrence across all emails.
Category Topics
Legal and Regulatory Affairs legal and contractual issues
legal department contacts
compliance and pipeline management
LNG financing opportunities
migration issues
BHP market inquiry
customer service inquiries
article reviews
Energy Markets and Operations California energy crisis
NYMEX website issues
energy market strategies
energy index management
accessing westpower desk
energy portfolio management
draft review process
document management issues
Administrative Coordination scheduling meetings
task or role reassignment
communication coordination
document review process
time zone conversion
recognition and support
variance methodology discussion
project collaboration tools
Corporate Strategy and Projects strategy session updates
financial data publication
Vince Kaminski project
document management issues
election concession parodies
fair trade opinions
Technology and Tools blackberry handheld devices
password security procedures
communication issues
Social and Interpersonal Communication personal relationships and communication
informal workplace conversations
miscommunication apologies
taste and acquired preferences
game interactions
Keneally’s social night
sailing lessons in Australia
wine retail pricing
socializing and drinks
congratulations and well wishes
family communications network
Miscellaneous and Culture horoscope and relationships
independence day plans
First, emails are preprocessed by removing punctuation, lemmatization, stopwords, and documents with less than 5 words. Then, we consider a transformers-based topic modeling method (Grootendorst, 2022) to extract latent topics from the Enron email dataset. In the procedure, texts are first embedded using pre-trained transformer ‘ll-MiniLM-L6-v2’, and then go through a dimension reduction step via uniform manifold approximation and projection (UMAP) (McInnes et al., 2018). The embeddings are clustered to identify topics, where the theme of each topic is extracted using cluster-based TF-IDF (Sparck Jones, 1972), and then fine-tuned using the GPT-4o Mini language model. This results in 47 well-defined and distinct topics (see Table 3). More details of this data processing procedure can be found in Section S8.
Our analysis focuses on the emails of 154 employees whose roles and departments are documented in the dataset. We begin by applying topic modeling to the full set of emails from all three stages to extract topics. Each email is then assigned one or more topics based on its content. To construct the network for each stage, we proceed as follows: for each pair of employees, we aggregate all emails between them in a given stage and construct a length-$K$ binary vector indicating the presence or absence of each topic in their communication. This process results in an undirected, multi-layer network for each stage with $n=154$ nodes (employees) and $K=47$ layers (topics). The presence of an edge in layer $k$ is denoted by $y_{ij}^{(k)}=y_{ji}^{(k)}=1$ if any email between users $i$ and $j$ is assigned topic $k$, and $y_{ij}^{(k)}=y_{ji}^{(k)}=0$ otherwise. This aggregation approach helps to reduce the sparsity of the network and facilitates comparison with other multi-layer network analysis methods. The edge densities of the resulting networks for stages 1, 2, and 3 are 0.23%, 0.22%, and 0.13%, respectively.